StepIV Shutdown Summer


This page contains the prioritised list of tasks for the Aug16 shutdown. The priorities are noted at end of each line. Priorities have been assigned on a scale of decreasing priority from 0 to 3. The specification of the priority categories are:
  • [0]: Critical; further operation will not take place before the task is complete;
  • [1]: Essential; task required to be complete by the end of the Aug16 shutdown;
  • [2]: Important; desirable that task be completed before the end of the Aug16 shutdown. Work on a priority [2] task must not displace work on tasks of priority [0] or [1];
  • [3]: Valuable; valuable that task is completed before the end of the Aug16 shutdown. Task can be completed later if work on the task during the shutdown would displace work on other tasks.

Task list:


Hall infrastructure

Conventional magnets
  • Cooling-water controls on roof1
  • Review existing water flows and temperature ranges (H.Nebrensky) [2]
Cooling cell
  • Hall probe improvement. [2]
    • Strain relief
    • Windows machines - MU plus AK


Daresbury Lab

Conventional Magnets (Here on 5/September)
  • Calibrate D1 & D2 current reading and update IOC software* (DL) 1-2 days [2]
Decay solenoid
  • Output current calibration verify read backs from front panel & GUI using independent DCCT, cross reference with input demand.
- 2 days
    • On hold until DS is up and working
  • Replace final battery cell [2]
Detectors, monitoring and control
  • Investigate archive data ‘dropouts’ [1] - being worked on by Adrian Oates
  • Increase resolution of PV measurement in archive viewer [2] : Stands
  • Increase resolution for cold mass and Helium temps (+1 digit) [2] : Stands



  • Channel
    • Set current/set current limit/set voltage limit. (in progress - model works, but needs implementation - also used for SS)
      • Sub-window for confirmation of any change. [0]
    • Buttons ( in progress - model works, but needs implementation)
      • Ramp [0] /zero [0] /pause [0]
      • Ramp enable [0]
      • Open close contractor [0] - tomorrow (DL team coming)
      • Sub-window for confirmation of any change [0]
  • Interlocks: Helium level/vac pressure (need Daresbury to work on it)
    • Require full interlock string before enable
    • He pressure sensor – reduce deadband [2]
      • To be done
  • Conventional Magnets
    • trips need to be in ALH. [1]: requires updating SM
  • Check spreadsheets with subsystem owners (limits/states etc.). [1]
  • Environment
    • Hall temperatures, monitoring + alarms. [2]
    • Neutron monitor alarms + NM fault alarm. [3]
  • SSH lockout [1] : Goes into Run Control. SB and PMH need to talk to define protocol.
    • Needs implementation
  • Target into ALH. [2]
    • Target IOC updated with Ed last week
    • Require spreadsheet for SM


Computing and Software

  • Fix DNS: servers give inconsistent results - DR [1] : awaits Paolo's return from holiday -Still broken-
  • Heartbeat vs. Nagios and alarms - Running: to be tested - PF. [1]
  • Check spare EMR board exists with correct Firmware (YK) [2] : Kingdom lost but for a cable. Board exist. Firmware exists. Cable doesn't.
    • What sort of cable is this - can we just buy one?
  • Ed will check with Yordan that automated disk management is implemented.
  • Get Target Depth, Target Delay, Total Beam Loss to read out (currently 0 or blank) (Ed/Pierrick) [1]
    • All this info pushed to EPICS. EPICS->CDB side needs work (Janusz)
  • Data Quality Flags #1383 (All, Janusz, Pierrick) [2] : Will not happen this user run
  • Reco: Data Quality flags #1585 (Durga, Pierrick) [2] : May happen this user run
  • Hardware token monitoring and alarms* #1860 H. Nebrensky [2]
    • Janusz has updated his Nagios probe to do the right alarm. Henry will speak to Paolo.
    • Need to respond to Nagios alarm by having somebody take action in MLCR. Not sure who - no CROC!


Updated by Franchini, Paolo over 7 years ago · 66 revisions